OpenStudy (anonymous):

If f(x) = -2x + 1, find f(-6). 12 13 -11 -7

OpenStudy (anonymous):

All what have you to do is to put -6 in place of x in the expression of f(x), can you have a try ?

OpenStudy (anonymous):

if given f(x) = -2x + 1 f(0) = -2*0 +1 f(-1) = -2* -1 +1 f( 10) = -2*10 +1 hopefully you see the pattern with functions

OpenStudy (anonymous):

f(-6)=-2(-6)+1 =12+1 =13